A researcher studying the nutritional value of a new candy places a 6.00 g sample of the candy inside a bomb calorimeter and combusts it in excess oxygen. The observed temperature increase is 2.23 ∘ C. If the heat capacity of the calorimeter is 37.20 kJ ⋅ K − 1 , how many nutritional Calories are there per gram of the candy?

Calories per gram of candy: _____Cal / g
Answer #1

Answer:-

Firstly total amount of heat released is calculated by using the simple formula and then calories per gram is calculated.

The answer is given in the image,

Answer a mass of candy=6.00g. Cecel = 37-2613-1-37-2k18-1 11=2-23°C Heat released - Cecel. AT = 31,2x2.23 = 82.96 KJ - 82-961